OpenStudy (anonymous):

Simplify. The variable represents a nonnegative real number. 3√98y^9

OpenStudy (anonymous):

\[3\sqrt{98y ^{9}}=3\sqrt{49*2*y ^{8}*y}=3*7*y ^{4} \sqrt{2y}=21y ^{4}\sqrt{2y}\]